tara read 19 science fiction and mystery novels in 6 months.she read 3 more science fiction novels than mystery novels. how many novels of each type did she read?
Answer by Maths68(1474)

Let
Science Fiction = s
Mystery Novels = m
Given
science fiction + Mystery Novels = 19
s+m=19........(1)
she read 3 more science fiction novels than mystery novels
s=m+3.........(2)

Plug in the value of s in (1)
s+m=19........(1)
m+3+m=19
2m+3=19
2m=19-3
2m=16
Divide by 2 both sides of the equation
2m/2=16/2
m=8
Put the value of m in (2)
s=m+3.........(2)
s=8+3
s=11
She reads
Science Fiction = s = 11
Mystery Novels = m = 8